# Measuring the Resolved Star Formation Main Sequence in TNG100: Fitting Technique Matters

### v1

@@ Abstract
The resolved star formation main sequence (rSFMS) and similar spatially resolved scaling relationships are frequently measured in both observed and simulated data. However, comparisons of these measurements are hindered by various differences between studies such as spatial resolution, sample selection criteria, and fitting technique. Using maps of $\Sigma_{*}$ and $\Sigma_{\rm SFR}$ derived using simulated galaxies from TNG100, we investigate the dependence of rSFMS measurements on spatial resolution, smoothing scale, and fitting technique. When the ridge line of the rSFMS is fit with a double linear function, we find that the slope of the rSFMS at low-$\Sigma_{*}$ is independent (within $2\sigma$) of spatial resolution and smoothing scale. We demonstrate that this method offers significant benefits over fitting the rSFMS with a single linear function using ordinary least squares.
@@
software: SciPy (Virtanen et al., 2020); ScaleRPy (McDonough, 2025)
@@ 1 Introduction
Scaling relationships are strong correlations between galaxy properties that arise due to the physical mechanisms that drive galaxy evolution. In particular, at low redshifts there exist relationships between the surface densities of gas mass, stellar mass ($\Sigma_{*}$), and star formation rate ($\Sigma_{\rm SFR}$) on sub-kiloparsec scales. These relationships are inherently linked to the evolution of galaxies, as they describe the formation of stellar populations. The relationship between $\Sigma_{*}$ and $\Sigma_{\rm SFR}$ is often called the resolved star formation main sequence (rSFMS; e.g., Cano-Díaz et al., 2016; Maragkoudakis et al., 2016; Abdurro’uf & Akiyama, 2017; Liu et al., 2018; Trayford & Schaye, 2019; Erroz-Ferrer et al., 2019; Medling et al., 2018; Lin et al., 2019; Bluck et al., 2020a; Pessa et al., 2021) and is the focus of this Letter.
@@
Scaling relationships play an important role in cosmological hydrodynamic simulations of galaxy evolution and unconstrained parameters can be calibrated to reproduce select relationships. For example, the IllustrisTNG simulations (Nelson et al., 2018; Pillepich et al., 2018; Springel et al., 2018; Naiman et al., 2018; Marinacci et al., 2018; Nelson et al., 2019) were calibrated to reproduce the stellar-to-halo mass, stellar mass – stellar size, and black hole – galaxy mass relations. Relationships that are not used to calibrate a simulation can, instead, be used to validate it. For example, in McDonough et al. (2023) we showed that the $100-$Mpc box of the IllustrisTNG suite (TNG100) broadly reproduced the rSFMS with a slope in agreement with that found by Bluck et al. (2020a) for MaNGA (Bundy et al., 2015) galaxies.
@@
As both simulations and observations improve in resolution and complexity, a need for accurate and unbiased quantitative comparisons arises. Unfortunately, differences in galaxy sample selection, spatial resolution, pre-selection of star-forming regions, star formation timescales, and techniques for measuring scaling relationships have made such comparisons challenging. The slopes of the rSFMS presented in the literature range from $\sim 0.5$ to $\sim 1$, with reported confidence intervals ranging from $\sim 0.0002-0.2$ (e.g., Cano-Díaz et al., 2016; Maragkoudakis et al., 2016; Abdurro’uf & Akiyama, 2017; Liu et al., 2018; Trayford & Schaye, 2019; Erroz-Ferrer et al., 2019; Bluck et al., 2020a; Hani et al., 2020; Pessa et al., 2021; McDonough et al., 2023). Using spatially resolved maps of simulated galaxies from the FIRE project (Wetzel et al., 2023), Hani et al. (2020) found that the rSFMS slope depends on the fitting method, star formation timescales, and spatial resolution.
@@
The rSFMS is likely a byproduct of two more fundamental relations (e.g., Baker et al., 2023): the molecular gas main sequence (i.e., $\Sigma_{H_{2}}-\Sigma_{*}$) and the Kennicutt-Schmidt relation (i.e., $\Sigma_{H_{2}}-\Sigma_{\rm SFR}$). However, the rSFMS remains useful as a proxy for the more fundamental relations due to the challenges associated with obtaining molecular gas measurements for large observational samples (i.e., the need for millimeter wavelengths). Even in simulation space, the rSFMS requires only the analysis of stellar particles, while investigations of the other relations would require analysis of both stellar and gas particles. Thus, a consistent measurement of the rSFMS slope is desirable for validation of simulation results and for probing the physics that drive star formation. Further, a consistent method for measuring the rSFMS relation should also be applicable to the more fundamental scaling relations.
@@
Many studies in the literature use some variation of ordinary least squares (OLS) to fit a line to their samples of spectral pixels (i.e., spaxels). Typically, some criteria are used to pre-select star-forming spaxels and a line is then fit to either the full sample or the median of the sample in bins of stellar mass. This method is not ideal because spaxels are not normally distributed about the rSFMS; as regions of galaxies quench, the SFR of spaxels will decrease while stellar mass remains roughly fixed. The introduction of pre-selection criteria for star-forming spaxels mitigates this issue to some degree, but is hard to apply consistently across studies, especially those using different SFR tracers.
@@
Renzini & Peng (2015) describe an objective definition for the global star-forming main sequence (SFMS; i.e., $M_{*}-{\rm SFR}$ relationship) based on the “ridge line” of data in 3D $M_{*}-{\rm SFR}$ plots. The ridge line method works by fitting a line to the mode of $\Sigma_{\rm SFR}$ in bins of $\Sigma_{*}$. This method is intended to ameliorate biases introduced by the presence of quenching spaxels without pre-selecting star-forming systems. The ridge line method has been adopted to measure the rSFMS in both observations (Abdurro’uf & Akiyama, 2017; Bluck et al., 2020a) and simulations (McDonough et al., 2023).
@@
In this letter, we investigate how different methodologies for obtaining maps of $\Sigma_{*}$ and $\Sigma_{\rm SFR}$ from galaxies in the TNG100 simulation affect the resulting rSFMS. We demonstrate that these effects are minimized when the slope is obtained by fitting a double linear function to the ridge line. In Section 2, we discuss the TNG simulations and our methodology for obtaining spatially-resolved data. In Section 3, we present our results for rSFMS fits obtained with: [i] different fitting techniques, [ii] spatially-resolved maps that have different resolutions, and [ii] different adaptive smoothing scales. In Section 4, we summarize our results and present our prescription for measuring unbiased, objective rSFMS slopes. To ensure the reproducibility of this work, the data and Jupyter notebook containing the analysis software is preserved on Zenodo: 10.5281/zenodo.15047581284 (catalog doi: 10.5281/zenodo.15047581284).
@@ 2 Data and Methods
Our galaxy sample is identical to that of McDonough et al. (2023) and McDonough et al. (2025), and was drawn from the TNG100 simulation. TNG100 used the IllustrisTNG model of galaxy formation, which adopted a Planck Collaboration (2016) cosmology. TNG100 has a comoving box length of $75h^{-1}\textrm{Mpc}$ and a target baryonic particle mass of $\sim 10^{6}h^{-1}M_{\odot}$. The IllustrisTNG model was calibrated such that at $z=0$ it reproduces the observed stellar mass function, the total gas mass of massive groups, and the following relations: stellar-to-halo mass, stellar mass to stellar size, and black hole to galaxy mass (Pillepich et al., 2018). Additionally, the model was calibrated to reproduce the shape of the cosmic star formation rate density as a function of redshift out to $z=10$.
@@
where $r$ is the radius a particle is smoothed over, $h_{sml}$ is the distance between a particle and its Nth nearest neighbor, and $q\equiv r/h_{sml}$. In McDonough et al. (2023), we presented our fiducial results for the rSFMS using 2D maps with a spatial resolution of $500h^{-1}{\rm pc}$ ($\sim 750$ pc) and a value of $N=64$ for the cubic spline kernel. Here, we obtain results from maps with different spatial resolutions and different values of N. Using $N=64$ we construct maps with spatial resolutions of $500$ pc, $\sim 750$ pc, $1$ kpc, and $\sim 1.5$ kpc. At a spatial resolution of $\sim 750$ pc, we also construct maps using $N=8$, $16$, $24$, $32$, and $64$. Pixels in 2D maps of galaxies from data obtained with integral field spectroscopy (IFS) are often referred to as spectral pixels or ‘spaxels.’ While the pixels in our 2D maps are not true spaxels, we refer to them as such for the sake of clarity.
@@
In Section 3.1, we present the rSFMS slope that we obtain using different methods to fit spatially resolved data. The rSFMS slope is often obtained from a fit to a single linear function (in logarithmic space). However, several studies have found that there is a point at which the global SFMS turns over (e.g., Renzini & Peng, 2015; Popesso et al., 2019; Mancini et al., 2019), and the bend also appears in the rSFMS (e.g., McDonough et al., 2023). Any fits made to a single linear function are either biased by the rSFMS turnover or must pre-select the regime where the rSFMS is purely linear.
@@
Although we mainly fit the ridge line to a double linear function to identify the rSFMS, we compare results for the rSFMS slope using variations of OLS fits in §3.1. For the non-OLS fits, we use the open-source ScaleRPy [^1] https://github.com/bryannemcd/ScaleRPy (McDonough, 2025) Python package to identify and fit the ridge line. In each bin of $\Sigma_{*}$, we use a Gaussian kernel density estimator (KDE; e.g., Scott, 1992) to identify the mode of $\Sigma_{\rm SFR}$. The ridge of our data is defined as the collection of points that correspond to the centers of the $\Sigma_{*}$ bins and the mode of the KDE of $\Sigma_{\rm SFR}$ in those bins. The ridge line and fits are computed over $10^{6}<\Sigma_{*}/[M_{\odot}{\rm kpc}^{2}]<10^{9}$, where there is sufficient data to obtain a KDE for each mass bin. The ridge line is fit with a double linear function parameterized as:
$\centering\log_{10}\Sigma_{\rm SFR,MS}=\begin{cases}\gamma\log_{10}\Sigma_{*}+y_{0}&\text{if }=\Sigma_{*}<\Sigma_{*}^{\rm to},\\ \gamma^{\prime}\log_{10}\Sigma_{*}+y_{0}^{\prime}&\text{if }\Sigma_{*}>\Sigma_{*}^{\rm to},\end{cases}\@add@centering$ (2)
@@
where $\gamma$ is the rSFMS slope, $y_{0}$ is the y-axis zero point, $\gamma^{\prime}$ is the slope of the high-mass end of the rSFMS, $\Sigma_{*}^{\rm to}$ is the value of $\Sigma_{*}$ where the main sequence turns over, and $y_{0}^{\prime}\equiv(\gamma-\gamma^{\prime})\log_{10}\Sigma_{*}^{\rm to}+y_{0}$.
@@
Figure 1: An example of the rSFMS from our fiducial maps, which demonstrates the ridge line and ordinary least squares fitting techniques. The 2D histogram shows the number of spaxels in bins of $\Sigma_{*}$ and $\Sigma_{\rm SFR}$. Yellow points indicate the ‘ridge’ or mode of the data in bins of $\Sigma_{*}$, and the black line shows the double linear fit to those ridges. The grey line shows the fit to a single linear function obtained using OLS and the full data set.
@@
In Figure 1, we compare the fits obtained with OLS (grey) and the ridge line (black) methods for the distribution of $\Sigma_{*}$-$\Sigma_{\rm SFR}$ from our fiducial maps. The yellow points indicate the ridges in each bin of $\Sigma_{*}$. For consistency with other investigations, we fit a single linear function with OLS (gray). However, the method we present here fits a double linear function to the ridge line (black) to account for the turnover in the rSFMS. Throughout, we use the $R^{2}$ statistic, also known as the coefficient of determination, to measure the goodness of our fits. A value of $R^{2}=1$ means that the model fits the data perfectly.

### v3

# Measuring Scaling Relations: Fitting Technique Matters

@@ Abstract
Scaling relationships, both integrated and spatially resolved, arise due to the physical processes that govern galaxy evolution and are frequently measured in both observed and simulated data. However, the accuracy and comparability of these measurements are hindered by various differences between studies such as spatial resolution, sample selection criteria, and fitting technique. Here, we compare variations of standard least squares techniques to the ridge line method for identifying spatially resolved scaling relations ($\Sigma_{*}-\Sigma_{\rm SFR}$, $\Sigma_{*}-\Sigma_{\rm gas}$, and $\Sigma_{\rm gas}-\Sigma_{\rm SFR}$) for TNG100 galaxies. We find that using the ridge line technique to fit these scaling relations with a double linear function results in significantly better fits than fitting with ordinary least squares. We further illustrate the utility of the ridge line technique with an investigation into the dependence of rSFMS measurements on spatial resolution and smoothing scale. Specifically, we find that the slope of the rSFMS at low-$\Sigma_{*}$ is independent (within $2\sigma$) of spatial resolution and smoothing scale. Finally, we discuss the need for a consistent re-analysis of resolved scaling relations in the literature and physically motivate adoption of the ridge line technique over other fitting methods.
@@
software: SciPy (Virtanen et al., 2020); ScaleRPy (McDonough, 2025); Plotly (Kruchten et al., 2024)
@@ 1 Introduction
Scaling relationships are strong correlations between galaxy properties that arise due to the physical mechanisms that drive galaxy evolution. As ever higher spatial resolutions have become possible, many relations previously identified between integrated (i.e., global) galaxy properties have been found to have spatially resolved analogs, down to $\sim 100$s of parsecs (e.g., Pessa et al., 2021). In particular, at low redshifts there exist relationships between the surface densities of gas mass ($\Sigma_{\rm gas}$), stellar mass ($\Sigma_{*}$), and star formation rate ($\Sigma_{\rm SFR}$) on sub-kiloparsec scales (see Sánchez, 2020, for a review). These relationships are inherently linked to the evolution of galaxies, as they describe the formation of stellar populations. The relationship between $\Sigma_{*}$ and $\Sigma_{\rm SFR}$ is often called the resolved star formation main sequence (rSFMS; e.g., Cano-Díaz et al., 2016; Maragkoudakis et al., 2016; Abdurro’uf & Akiyama, 2017; Liu et al., 2018; Trayford & Schaye, 2019; Erroz-Ferrer et al., 2019; Medling et al., 2018; Lin et al., 2019; Bluck et al., 2020a; Pessa et al., 2021; Barrera-Ballesteros et al., 2021) and is the main focus of this Letter. However, we also present the resolved gas main sequence (rGMS, i.e., $\Sigma_{*}-\Sigma_{\rm gas}$ relation) and resolved Kennicutt-Schmidt law (rKS, i.e., $\Sigma_{\rm gas}-\Sigma_{\rm SFR}$ relation) for simulated galaxies.
@@
Scaling relationships play an important role in cosmological hydrodynamic simulations of galaxy evolution, since unconstrained parameters can be calibrated to reproduce select relationships. For example, the IllustrisTNG simulations (Nelson et al., 2018; Pillepich et al., 2018; Springel et al., 2018; Naiman et al., 2018; Marinacci et al., 2018; Nelson et al., 2019) were calibrated to reproduce the stellar-to-halo mass, stellar mass – stellar size, and black hole – galaxy mass relations. Relationships that are not used to calibrate a simulation can, instead, be used to validate it. For example, in McDonough et al. (2023) we showed that the $100-$Mpc box of the IllustrisTNG suite (TNG100) broadly reproduced the rSFMS with a slope in agreement with that found by Bluck et al. (2020a) for MaNGA (Bundy et al., 2015) galaxies.
In general, scaling relations have broad utility in extragalactic astrophysics. For observational measurements that are challenging to make, scaling relations can be used to identify appropriate proxies, e.g., inferring black hole mass from central velocity dispersion (e.g., Gebhardt et al., 2000). Semi-analytic models often incorporate scaling relations, for example: populating simulated dark matter halos with galaxies given the halo-to-stellar mass relation (see De Lucia, 2019, for a review). Further, scaling relations arise due to physical processes, and therefore their shapes encode physically significant information. For example, Schmidt (1959) demonstrated a power-law relationship between SFR and the surface density of gas. Kennicutt (1989) and Kennicutt (1998) expanded on this work, finding that the SFR was also regulated by the total local density which sets the conditions for gravitational collapse. Given that the timescale for gravitational instabilities to grow in a parcel of gas will scale with local density, a star formation law should follow the form: $\rho_{\rm SFR}\propto\rho_{\rm gas}^{1.5}$ above some threshold density (Kennicutt, 1998). More recently, Shi et al. (2011) proposed an extension of the rKS law with $\Sigma_{\rm SFR}$ being dependent on both $\Sigma_{\rm gas}$ and $\Sigma_{*}$. Many works have found that the rKS holds only for molecular gas (e.g., Kennicutt et al., 2007). A full review of theoretical predictions for the rKS and similar relations is beyond the scope of this Letter, but in general the slopes of these power-law relations are predicted by theory and can be constrained with observational measurements.
@@
As both simulations and observations improve in resolution and complexity, a need for accurate and unbiased quantitative comparisons of scaling relations arises if we are to constrain theory further. Of particular interest is whether the physical properties that drive spatially resolved scaling relations are redshift-dependent, as has been found for some global relations, e.g., the global SFMS (Speagle et al., 2014). Unfortunately, differences in galaxy sample selection, spatial resolution, and techniques for measuring scaling relationships have made such comparisons challenging. Measurements of the rSFMS and rKS can be further affected by the criteria for pre-selection of star-forming regions and the indicator used to measure SFR. The slopes of the rSFMS presented in the literature range from $\sim 0.5$ to $\sim 1$, with reported confidence intervals ranging from $\sim 0.0002-0.2$ (e.g., Cano-Díaz et al., 2016; Maragkoudakis et al., 2016; Abdurro’uf & Akiyama, 2017; Liu et al., 2018; Trayford & Schaye, 2019; Erroz-Ferrer et al., 2019; Bluck et al., 2020a; Hani et al., 2020; Pessa et al., 2021; McDonough et al., 2023). Using spatially resolved maps of simulated galaxies from the FIRE project (Wetzel et al., 2023), Hani et al. (2020) found that the rSFMS slope depends on the fitting method, star formation timescales, and spatial resolution.
@@
The rSFMS may be a byproduct of two more fundamental relations (e.g., Baker et al., 2023): the molecular gas main sequence (i.e., $\Sigma_{H_{2}}-\Sigma_{*}$) and the molecular rKS law (i.e., $\Sigma_{H_{2}}-\Sigma_{\rm SFR}$). However, the rSFMS remains useful due to the challenges associated with obtaining molecular gas measurements for large observational samples (i.e., the need for millimeter wavelengths). Even in simulation space, the rSFMS requires only the analysis of stellar particles, while investigations of the other relations requires analysis of both stellar and gas particles. Thus, a consistent measurement of the rSFMS slope and other scaling relations is desirable for validation of simulation results and for probing the physics that drive star formation.
@@
Many studies in the literature use some variation of least squares to fit a power law to their samples of spectral pixels (i.e., spaxels) or simulated analogs, usually done in logarithmic space to simplify the problem. Typically, some criteria are used to pre-select star-forming spaxels and a line is then fit to either the full sample or the median of the sample in bins of stellar mass. Mathematically, a least squares fit to a full sample returns a line that traces the conditional mean (e.g., E$[\Sigma_{\rm SFR}|\Sigma_{*}]$). This method is not ideal because spaxels are not normally distributed about the rSFMS; as regions of galaxies quench, the SFR of spaxels will decrease while stellar mass remains roughly fixed. Fitting a line to the median of the data (i.e., identifying a linear approximation to the conditional median) limits the influence of quenched regions, although it is still biased by non-normal distributions. The introduction of pre-selection criteria for star-forming spaxels may limit the data to the regime where it is normally distributed, but is hard to apply consistently across studies, especially those using different SFR tracers. More broadly, any scaling relations fit with least squares can be biased by the presence of outliers, non-normal scatter and other limitations on the range of data covered (e.g., magnitude- or mass- limited samples).
@@
Renzini & Peng (2015) describe an objective definition for the global star-forming main sequence (SFMS; i.e., $M_{*}-{\rm SFR}$ relationship) based on the “ridge line” of data in 3D $M_{*}-{\rm SFR}$ plots. The ridge line method works by fitting a line to the mode of $\Sigma_{\rm SFR}$ in bins of $\Sigma_{*}$ (i.e., the conditional mode). This method is intended to ameliorate biases introduced by the presence of quenching spaxels without pre-selecting star-forming systems. Mathematically, the ridge line accomplishes this because it returns a function that traces the maximum of the conditional probability distribution function (e.g., max[P$(\Sigma_{\rm SFR}|\Sigma_{*})$]). The ridge line method has been adopted to measure the rSFMS in both observations (e.g., Abdurro’uf & Akiyama, 2017) and simulations (McDonough et al., 2023).
@@
In this Letter, we investigate how different methodologies for obtaining maps of $\Sigma_{*}$ and $\Sigma_{\rm SFR}$ from galaxies in the TNG100 simulation affect the resulting rSFMS. We demonstrate that these effects are minimized when the slope is obtained by fitting a double linear function to the ridge line. Our ridge line method builds on the approach proposed by Renzini & Peng (2015), and we hope to further facilitate the adoption of this technique by releasing the open-source Python package ScaleRPy (McDonough, 2025). In Section 2, we discuss the TNG simulations and our methodology for obtaining spatially-resolved data. We also present the rSFMS, rGMS, and rKS for TNG100 galaxies. In Section 3.1, we present our results for rSFMS, rKS, and rGMS fits obtained with different fitting techniques. In Section 3.2 and 3.3, we compare ridge line fits of the rSFMS from galaxy maps generated at different spatial resolutions and with different adaptive smoothing scales, respectively. In Section 4, we summarize our results and present our prescription for measuring unbiased, objective slopes of scaling relations. To ensure the reproducibility of this work, the data and Jupyter notebook containing the analysis software is preserved on Zenodo: 10.5281/zenodo.15047580 (catalog doi: 10.5281/zenodo.15047580). The code repository for ScaleRPy is hosted on Github [^1] ScaleRPy codebase: https://github.com/bryannemcd/ScaleRPy. and the specific version of ScaleRPy used for this Letter is permanently archived on Zenodo (McDonough, 2025).
@@ 2 Data and Methods
Our galaxy sample is identical to that of McDonough et al. (2023) and McDonough et al. (2025), and was drawn from the TNG100 simulation. TNG100 used the IllustrisTNG model of galaxy formation, which adopted a Planck Collaboration (2016) cosmology. TNG100 has a comoving box length of $75h^{-1}\textrm{Mpc}$ and a target baryonic particle mass of $\sim 10^{6}h^{-1}M_{\odot}$. The IllustrisTNG model was calibrated such that at $z=0$ it reproduces the observed stellar mass function, the total gas mass of massive groups, and the following scaling relations: stellar-to-halo mass, stellar mass to stellar size, and black hole to galaxy mass (Pillepich et al., 2018). Additionally, the model was calibrated to reproduce the shape of the cosmic star formation rate density as a function of redshift out to $z=10$. The TNG simulations use a star formation prescription (Springel & Hernquist, 2003) with free parameters tuned to reproduce the theoretical rKS slope of $1.5$ proposed by Kennicutt (1998).
@@
where $r$ is the radius a particle is smoothed over, $h_{sml}$ is the distance between a particle and its Nth nearest neighbor of the same particle type, and $q\equiv r/h_{sml}$. For this work, we have also generated maps of total gas mass using the same kernel.
We choose as our fiducial parameters those used in McDonough et al. (2023), which generated 2D maps with a spatial resolution of $500h^{-1}{\rm pc}$ ($\sim 750$ pc) and a value of $N=64$ for the cubic spline kernel. The fiducial spatial resolution was selected to be approximately comparable to the resolution of MaNGA data used by Bluck et al. (2020a) and (Bluck et al., 2020b). The fiducial value of $N$ was selected to improve the relative coverage of star-forming regions. Additional maps of $\Sigma_{*}$ and $\Sigma_{\rm SFR}$ were generated at different spatial resolutions and different values of $N$ to explore the effect of these parameters on the resulting fit. Using $N=64$ we construct maps with spatial resolutions of $500$ pc, $\sim 750$ pc, $1$ kpc, and $\sim 1.5$ kpc. At a spatial resolution of $\sim 750$ pc, we also construct maps using $N=8$, $16$, $24$, $32$, and $64$. The values of $N$ that we explore here were inspired by the range of $N$ over which Rodriguez-Gomez et al. (2019) found their morphological measurements of TNG100 galaxies to be insensitive. We note that this parameter does not have a direct analog in observed data, but the choice of $N$ for simulated galaxies impacts the spatial extent of star-forming regions and the patchiness of stellar mass maps. Pixels in 2D maps of galaxies from data obtained with integral field spectroscopy (IFS) are often referred to as spectral pixels or ‘spaxels.’ While the pixels in our 2D maps are not true spaxels, we refer to them as such for the sake of clarity. We limit our analysis to spaxels with $\Sigma_{\rm SFR}>10^{6}M_{\odot}{\rm kpc}^{2}$, consistent with Bluck et al. (2020b) and McDonough et al. (2023). For the rKS and rGMS relations, we further limit our analysis to spaxels with $\Sigma_{\rm gas}>10^{6}M_{\odot}{\rm kpc}^{2}$, the approximate observational limit of recent high-resolution surveys of nearby galaxies (e.g., Pessa et al., 2021).
@@
In Section 3.1, we present the rSFMS, rKS, and rGMS slopes that we obtain using five different methods to fit spatially resolved data: [1] ridge line: our fiducial approach described below; [2] OLS: ordinary least squares applied to the spaxel data in $\log_{10}$ space; [3] LS-1PL: non-linear least squares to fit the spaxel data to a single power law; [4] LS-2PL: non-linear least squares to fit the spaxel data to a double power law; and [5] LS-2PL-log: non-linear least squares to fit the spaxel data to a double linear in $\log_{10}$ space. While many studies have fit the spatially resolved scaling relations with a single power law, several studies have found that there is a point at which the global SFMS turns over (e.g., Renzini & Peng, 2015; Popesso et al., 2019; Mancini et al., 2019), and the bend also appears in the rSFMS (e.g., Abdurro’uf & Akiyama, 2017; McDonough et al., 2023). A bend in the rKS is also expected at the density threshold for star formation to occur (Kennicutt, 1989). Any rSFMS or rKS fits made to a single power law function are thus either biased by these transitions or must pre-select the regime where the relationship is purely linear in logarithmic space. For this reason, methods [1], [4], and [5] fit to the functional form of a double power law, and we compare this to a single power law with methods [2] and [3]. We further demonstrate the difference between fitting in linear space-- methods [2] and [4]--and logarithmic space-- methods[1], [3], and [5]. As we will show, this difference can be significant.
@@
We use the open-source ScaleRPy Python package to identify and fit the ridge line described by Renzini & Peng (2015). In each bin of $\log_{10}\Sigma_{*}$, we use a Gaussian kernel density estimator (KDE; e.g., Scott, 1992) to identify the mode of $\log_{10}\Sigma_{\rm SFR}$. The ridge of our data is defined as the collection of points that correspond to the centers of the $\log_{10}\Sigma_{*}$ bins and the mode of the KDE of $\log_{10}\Sigma_{\rm SFR}$ in those bins. Unlike a mean or median, the mode of a data set in linear space will be identical to the mode of a data set in $\log_{10}$ space. We choose to perform the ridge line fit in $\log_{10}$ space because the KDE software we use assumes a fixed bandwidth, which works poorly for data that span many orders of magnitude. We have tested that fitting the identified ridge points with least squares in linear space does not significantly affect our results. Uncertainties on the ridge points are estimated as the KDE bandwidth for a given bin of $\Sigma_{*}$ and determined via Scott’s rule (Scott, 1979), which considers the number of data points and their standard deviation from a normal distribution. The rSFMS and rGMS ridge line and fits are computed over $10^{6}<\Sigma_{*}/[M_{\odot}{\rm kpc}^{2}]<10^{9}$, where there is sufficient data to obtain a KDE for each mass bin in the rSFMS. The ridge line is fit with a double power law parameterized as:
$\centering\log_{10}\Sigma_{\rm SFR,MS}=\begin{cases}\gamma\log_{10}\Sigma_{*}+y_{0}&\text{if }\Sigma_{*}<\Sigma_{*}^{\rm to},\\ \gamma^{\prime}\log_{10}\Sigma_{*}+y_{0}^{\prime}&\text{if }\Sigma_{*}>\Sigma_{*}^{\rm to},\end{cases}\@add@centering$ (2)
@@
where $\gamma$ is the rSFMS slope, $y_{0}$ is the y-axis zero point, $\gamma^{\prime}$ is the slope of the high-mass end of the rSFMS, $\Sigma_{*}^{\rm to}$ is the value of $\Sigma_{*}$ where the main sequence turns over, and $y_{0}^{\prime}\equiv(\gamma-\gamma^{\prime})\log_{10}\Sigma_{*}^{\rm to}+y_{0}$. A similar parameterization is used to fit the rKS and the rGMS.
(a) rSFMS
(b) rKS
(c) rGMS
@@
Figure 1: Panels a-c: The rSFMS (a), rKS (b), and rGMS (c) scaling relations from our fiducial maps, which demonstrate the ridge line and ordinary least squares fitting techniques. The 2D histogram shows the number of spaxels (i.e., samples) in bins of values on the x and y axes. Yellow points indicate the ‘ridge’ or conditional mode of the data, and the black line shows the double linear fit to those ridges. Uncertainties on the ridge points are estimated as the bandwidth of the KDE obtained with Scott’s rule (Scott, 1979). The gray lines show least squares fits to a single power law (dashed and dot-dashed lines) and to a double power law (dotted and solid lines) using the full data set in linear space (dot-dashed and solid) and in $\log_{10}$ space (dashed and dotted). Panel d: A $3$D visualization of the relationship between $\Sigma_{*}$, $\Sigma_{\rm gas}$, and $\Sigma_{\rm SFR}$. Iso-density contours are drawn at the $50^{\rm th}$, $70^{\rm th}$, $85^{\rm th}$, $95^{\rm th}$, $99.9^{\rm th}$ percentiles. In the html version of the published Letter, this panel is available as an interactive figure, where the $3$D visualization can be rotated.
@@
In Figure 1, we compare the fits obtained with least squares (gray) and the ridge line (black) methods for the distributions of $\Sigma_{*}-\Sigma_{\rm SFR}$ (Figure 1a), $\Sigma_{\rm gas}-\Sigma_{\rm SFR}$ (Figure 1b), and $\Sigma_{*}-\Sigma_{\rm gas}$ (Figure 1c) from our fiducial maps. The yellow points indicate the ridges in each bin of $\Sigma_{*}$. The dashed lines are the fits obtained with OLS, dot-dashed lines are fits obtained with LS-1PL, dotted lines are fits obtained with LS-2PL-log, and the solid gray lines are fits obtained with LS-2PL. In Figure 1d, we visualize the distribution of spaxel data in 3D space with iso-density contours, to highlight the relationship between all three variables. Throughout, we use the $R^{2}$ statistic, also known as the coefficient of determination, to measure the goodness of our fits. A value of $R^{2}=1$ means that the model fits the data perfectly. Comparison of the $R^{2}$ statistic for ridge line and least squares fits are not necessarily straightforward, since the ridge line technique fits to far fewer points. Thus, the variation of data from least squares fits is naturally greater. However, this is one of the advantages of the ridge line technique: it first identifies what it is that should be fit (the maximum of the conditional probability density function) before fitting a relation.